OpenStudy (anonymous):

Calculate the area of triangle ABC with altitude CD, given A (6, 0), B (1, 5), C (2, 0), and D (4, 2) Select one: a. 5 square units b. 8 square units c. 10 square units d. 13 square units

OpenStudy (anonymous):

@Abmon98

OpenStudy (anonymous):

A = (1/2) BASE * HEIGHT

OpenStudy (anonymous):

What's the base and what's the height

OpenStudy (abmon98):

((1-6)^2+(5-0)^2)^1/2=5(2)^1/2 ((2-1)^2+(5-0)^2)^1/2=(26)^1/2 ((2-6)^2+(0-0)^2)^1/2=4 ((26)^1/2*4)/2=10 square units